OpenStudy (anonymous):

find the amount decrease and percent decrease Orginal Amount:80 New Amount: 60 Amount of Decrease?? Percent Decrease??

OpenStudy (turingtest):

amount: 80-60=20 percent: (20/80)x100=25 percent need more explanation?

OpenStudy (anonymous):

how did you get 25%?

OpenStudy (anonymous):

\[\frac{20}{80}=\frac{1}{4}=0.25=25\%\]